What is the simple method to findout the Integration of cosx divided by 2+sinx

Grade:12

2 Answers

Aaditya Sharma
7 years ago
It can be easily done by thr method of ‘u-substitution’.      
I=\int \frac{cosx}{2+sinx} dx
Put u=sinx
I=\int \frac{1}{2+u} du
 
I= In |u+2|+C
I= In |sinx+2|+C
Hope it helps.
rangu meghana
7 years ago
consider sinx as t 
sinx=t
differentiate on bothsides
cosx=dt\dx    => cosxdx=dt ,sub it in question
we get ∫dt/(2+t)  = log(2+t) =log(2+sinx)